OpenStudy (anonymous):

SAT Question - Help? The Hungry Ghosts are a rock band that can play 7 different songs. FOr an upcoming show, the Hungry Ghosts have time to play only 4 songs in their set. If a set consists of 4 songs in a specific order, how many different sets do the Hungry Ghosts have the option of playing at their upcoming show. Answer : 840 How do I do this?

OpenStudy (kropot72):

Have you studied permutations?

OpenStudy (anonymous):

Sort of? I know it's something order dependent.

OpenStudy (kropot72):

The n umber of permutations of n different things taken r at a time is: \[\large P(n, r)=\frac{n!}{(n-r)!}=\frac{7\times6\times5\times4\times3\times2\times1}{3\times2\times1}=you\ can\ calculate\]

OpenStudy (anonymous):

@kropot72 How about the formula for combinations? Does it look similar to that? And thanks for the formula much easier than how my teacher explained it.

OpenStudy (anonymous):

I get the top part of n! but can you explain what r is?

OpenStudy (kropot72):

The formula that I posted for permutations is 'without repetitions'. With repetitions the number of permutations is: \[\large n ^{r}\] The number of combinations of n different things taken r at a time is: \[\large C(n, r)=\frac{n!}{r!(n-r)!}\] r is the number of the n different things taken at a time. In the question that you posted, r is the number of songs in a set which is 4.

OpenStudy (anonymous):

Oh okay so it was 7-4 so find the r. So is that how you usually find r? Haha sorry if this is a dumb question.

OpenStudy (kropot72):

In the question, n = 7 being the total number of different songs. r = 4 being the number of songs taken at a time: \[\large n-r=7-4=3\]

OpenStudy (anonymous):

Thank you very much